OpenStudy (anonymous):

is 2^3 bigger or smaller than 3^4?

OpenStudy (hba):

2^3=2*2*2= 3^4=3*3*3*3= Now you decide :)

OpenStudy (anonymous):

3^4 is bigger?

OpenStudy (truehawaiian):

Yeah...

OpenStudy (anonymous):

ok thanks